OpenStudy (anonymous):

What is the solution to the proportion? 6/15 = x/20 A. x = 8 B. x = 2.5 C. x = 4.5 D. x = 12

OpenStudy (anonymous):

x = 6*20/15 = 8 (what i think) @NerdyMackenzie

OpenStudy (anonymous):

@NerdyMackenzie

OpenStudy (radar):

Nothing wrong with those kind of thoughts.

OpenStudy (anonymous):

You got it right :)

OpenStudy (anonymous):

woop woop :DDDDDDDDDDDDDDDDD

OpenStudy (anonymous):

i would give you a medal @radar but i already gave one out :L

OpenStudy (anonymous):

what is the solution to the problem ? m/18 = 15/12

OpenStudy (anonymous):

cross multiply 12m = 18*15 m = 18*15/12 =

OpenStudy (anonymous):

\(\bbox [ 1pt, #66f ,border: 17px solid #000 ] {\Huge\sf\color{white} {idk~this~one~~:/ }}\)

OpenStudy (anonymous):

22.5?

OpenStudy (anonymous):

Yes you got it right again :)

OpenStudy (anonymous):

\(\bbox [ 1pt, #66f ,border: 17px solid #000 ] {\Huge\sf\color{white} {22.5? }}\)

OpenStudy (anonymous):

lol im horrible at math =P

OpenStudy (anonymous):

but smart when i comes to xbox =P

OpenStudy (anonymous):

i know every gun on call of duty, but i cant remember how to divide some times =D

OpenStudy (anonymous):

brb gotta reload

OpenStudy (anonymous):

bck

OpenStudy (anonymous):

Practice makes perfect. But if your good in other topics just focus on those topics.

OpenStudy (anonymous):

Ten pencils cost $1.50. How much does it cost to buy 4 pencils? A. $0.60 B. $0.75 C. $3.75 D. $6.00

OpenStudy (anonymous):

10/ 1.50 then 15/4 = 0.60.....

OpenStudy (anonymous):

1.50/10=0.15 15*4=0.06

OpenStudy (anonymous):

idk if im right , my mom helped me lol

OpenStudy (anonymous):

I mean 0.60 so you go it right,

OpenStudy (anonymous):

wow, my mom sucks at math lol

OpenStudy (anonymous):

and no, im not a nerd T-T

OpenStudy (anonymous):

Do you need help in anything else?

OpenStudy (anonymous):

nope thanks alot :D

OpenStudy (anonymous):

your welcome :)